# Parabolic problems whose Fujita critical exponent is not given by scaling

### v1

@@ 1. Introduction
For further details on these fundamental properties of Riesz potentials, we refer the reader to [19, Section I.1], [28, Section V.1.1].
In addition, as $\alpha\to 0$, the Riesz potential $I_{\alpha}$ converges to the Dirac delta $\delta$ in the vague sense [19, p. 46]. On the other hand, as $\alpha\to n$, one has
@@
The definition of Riesz potentials can be extended from $\alpha\in(0,n)$ to arbitrary complex $\alpha$ with $\mathrm{Re}(\alpha)>0$ and $(\alpha-n)/2\notin\mathbb{N}\cup\{0\}$. In this more general setting, the convolution $I_{\alpha}f$ must be interpreted in the sense of distributions, see for instance [19] or [26, Chapter 2]. Throughout this paper, we restrict to the case $\alpha\in(0,n)$, where the convolution $I_{\alpha}f$ is understood in the classical Lebesgue integral sense.
@@ 1.1.3. Parabolic problems with convolution nonlinearities
then problem (1.1) has no nontrivial weak solutions. Here, $B_{R}$ stands for the closed ball centered at O with radius $R.$ However, the case $p>1+(2+\alpha)/(n-\alpha)$ was not explored in that paper, and the authors conjectured that (see [23, Remark 7]) when $p>1+(2+\alpha)/(n-\alpha),$ there exists a global solution corresponding to initial data $u_{0},$ which satisfies
@@ 1.2. Motivation
$p_{\mathrm{Fuj}}(n,\alpha,\beta)=1+\frac{2+\alpha}{n-\alpha}.$
@@ 1.3. Main results
As shown in Theorem 1.2, the local existence of a solution to problem (1.1) requires that $p>n/(n-\alpha)$. This restriction arises from the nonlocal nature of the nonlinear source term. Consequently, in the case $1<p\leq n/(n-\alpha),$ whether the local solution exists or not remains an open question.
@@
    $p_{\mathrm{Fuj}}(n,\beta,\alpha):=1+\frac{\beta+\alpha}{n-\alpha}.$
@@ 2. Preliminaries
The following Hardy-Littlewood-Sobolev is a particular case of [20, Theorem 4.3].
@@
[6, Lemma 2.4] Let $s\in(0,1]$, and $\varphi$ be a smooth function satisfying $\partial_{x}^{2}\varphi\in L^{\infty}(\mathbb{R}^{n})$. For any $R>0$, let $\varphi_{R}$ be a function defined by
@@
Applying [3, Lemma 4.1.5], using the continuity of the semigroup $S_{\beta}(t)$, we conclude that $u\in C([0,T],L^{s}(\mathbb{R}^{n}))$.
@@ 4. Blow-up
where $\mathcal{B}=\left\{x\in\mathbb{R}^{n};\,\,|x|\leq R\right\}$. Let us first derive an estimate for $I_{1}$. Using Hölder’s inequality together with Ju’s inequality $(-\Delta)^{\beta/2}\left(\varphi_{R}^{\ell}\right)\leq\ell\varphi_{R}^{\ell-1}(-\Delta)^{\beta/2}\varphi_{R}$ (see e.g. [9, Appendix]), we have
@@ Acknowledgments
Ahmad Fino is supported by the Research Group Unit, College of Engineering and Technology, American University of the Middle East. Berikbol T. Torebek is supported by the Science Committee of the Ministry of Education and Science of the Republic of Kazakhstan (Grant No. AP23483960)

### v2

@@ Abstract.
Corresponding author: Berikbol T. Torebek, email: torebek@math.kz
@@ 1. Introduction
For further details on these fundamental properties of Riesz potentials, we refer the reader to [20, Section I.1], [29, Section V.1.1].
In addition, as $\alpha\to 0$, the Riesz potential $I_{\alpha}$ converges to the Dirac delta $\delta$ in the vague sense [20, p. 46]. On the other hand, as $\alpha\to n$, one has
@@
The definition of Riesz potentials can be extended from $\alpha\in(0,n)$ to arbitrary complex $\alpha$ with $\mathrm{Re}(\alpha)>0$ and $(\alpha-n)/2\notin\mathbb{N}\cup\{0\}$. In this more general setting, the convolution $I_{\alpha}f$ must be interpreted in the sense of distributions, see for instance [20] or [27, Chapter 2]. Throughout this paper, we restrict to the case $\alpha\in(0,n)$, where the convolution $I_{\alpha}f$ is understood in the classical Lebesgue integral sense.
@@ 1.1.3. Parabolic problems with convolution nonlinearities
then problem (1.1) has no nontrivial weak solutions. Here, $B_{R}$ stands for the closed ball centered at O with radius $R.$ However, the case
$p>1+(2+\alpha)/(n-\alpha)$
was not explored in that paper, and the authors conjectured that (see [24, Remark 7]) when $p>1+(2+\alpha)/(n-\alpha),$ there exists a global solution corresponding to initial data $u_{0},$ which satisfies
@@ 1.2. Motivation
$p_{\mathrm{Fuj}}(n,2,\alpha)=1+\frac{2+\alpha}{n-\alpha}.$
@@ 1.3. Main results
As shown in Theorem 1.2, the local existence of a solution to problem (1.1) requires that $p>n/(n-\alpha)$. This restriction arises from the nonlocal nature of the nonlinear source term and, in particular, from the use of Lemma 2.2 combined with Hölder’s inequality, and seems to be of a technical nature. Consequently, in the case $1<p\leq n/(n-\alpha),$ whether the local solution exists or not remains an open question.
@@
    $p_{\mathrm{Fuj}}(n,\beta,\alpha)=1+\frac{\beta+\alpha}{n-\alpha}.$
@@
  • $(\mathrm{e})$
    To better understand the originality of the critical exponent $p_{\mathrm{Fuj}}(n,\beta,\alpha)$, let us note that the steady-state solution $v$ of problem (1.1) can formally satisfy
    $(-\Delta)^{\frac{\alpha+\beta}{2}}v=|v|^{p}$
    where we have used the fact that the Riesz potential $I_{\alpha}$ can be regarded as the inverse of the fractional Laplacian $(-\Delta)^{\alpha/2}$. By setting $\varphi(x)=|x|^{-(n-\alpha)}$ for $x\in\mathbb{R}^{n}\setminus\{0\}$ and applying the estimate of Bonforte and Vázquez [3]:
    $(-\Delta)^{\frac{\alpha+\beta}{2}}\varphi\leq C|x|^{-(n-\beta)},$
    one can see that $\varphi$ is a steady-state subsolution of problem (1.1) if
    $(n-\alpha)p=n+\beta\qquad\hbox{that is}\qquad p=p_{\mathrm{Fuj}}(n,\beta,\alpha).$
@@
Typical examples of $\mathcal{K}$ are the constant functions as well as
$\mathcal{K}(r)=r^{-(n-\alpha)},\,\,\alpha\in(0,n)\quad\hbox{or}\quad\mathcal{K}(r)=r^{-(n-\alpha)}\log^{\beta}(1+r),\,\,\alpha\in(0,n),\,\beta\in\mathbb{R},\,\beta+\alpha>0.$
@@
  • d)
    For the sake of clarity, we emphasize the distinction between the finite-time blow-up of mild solutions and the nonexistence of global weak solutions. Finite-time blow-up of mild solutions means that a solution exists on a maximal time interval $0,T_{\max})$ with $0<T_{\max}<\infty$, but the norm of the solution, in an appropriate functional space, becomes unbounded as $t\rightarrow T_{\max}$. In contrast, the nonexistence of global weak solutions refers to the fact that no solution, even in a weak sense, can be defined for all $t\geq 0$ under the given assumptions.
@@ 2. Preliminaries
The following Hardy-Littlewood-Sobolev is a particular case of [21, Theorem 4.3].
@@
[7, Lemma 2.4] Let $s\in(0,1]$, and $\varphi$ be a smooth function satisfying $\partial_{x}^{2}\varphi\in L^{\infty}(\mathbb{R}^{n})$. For any $R>0$, let $\varphi_{R}$ be a function defined by
@@
Applying [4, Lemma 4.1.5], using the continuity of the semigroup $S_{\beta}(t)$, we conclude that $u\in C([0,T],L^{s}(\mathbb{R}^{n}))$.
@@ 4. Blow-up
where $\mathcal{B}=\left\{x\in\mathbb{R}^{n};\,\,|x|\leq R\right\}$. Let us first derive an estimate for $I_{1}$. Using Hölder’s inequality together with Ju’s inequality $(-\Delta)^{\beta/2}\left(\varphi_{R}^{\ell}\right)\leq\ell\varphi_{R}^{\ell-1}(-\Delta)^{\beta/2}\varphi_{R}$ (see e.g. [10, Appendix]), we have
@@ Acknowledgments
Ahmad Fino is supported by the Research Group Unit, College of Engineering and Technology, American University of the Middle East. Berikbol T. Torebek is supported by the Science Committee of the Ministry of Education and Science of the Republic of Kazakhstan (Grant No. AP26195417)